### v2

# Energy Transport Among Highly-Polarized Atoms

Abstract
A static electric field of a few V/cm shifts the energy levels of ultracold Rydberg atoms in a magneto-optical trap. For a given principle quantum number, most of the energy levels are nearly degenerate at zero field and fan out with increasing field to form a manifold. We excite Rydberg atoms to energy levels near the center of the manifold, where the spacing is nearly harmonic, and allow them to exchange energy via resonant dipole-dipole interactions. We measure the time evolution as energy spreads away from the center of the manifold, which reveals that the system fails to thermalize for long interaction times. A computational model that includes only a few essential features of the system qualitatively agrees with this result.
Harmonic ladders of energy levels are ubiquitous in nature. Beyond the paradigmatic quantum harmonic oscillator, energy ladders are the basis for many phenomena in atomic, chemical, and condensed matter systems. They describe the motion of trapped ions [1, 2, 3, 4], arise from molecular and lattice vibrations [5, 6, 7], and emerge as Floquet states in strongly driven systems [8, 9, 10]. Landau levels can contribute to quantum Hall phenomena [11, 12], and have been recently observed in a variety of solid-state systems [13, 14, 15], including skyrmions [16, 17]. These energy structures lead to interesting studies of nonequilibrium dynamics, topological phases, and energy transport [18, 19, 20].
Harmonic ladders give rise to a range of modern applications, such as trapped-ion qubits for quantum information [21], or in quantum control of molecular disassociation through vibrational ladder climbing [22, 23, 24]. Manipulation of confinement in quantum dots has enabled tuning of optical and electrical properties for lasing, optoelectronic devices, and biological applications [25, 26, 27, 28, 29, 30, 31, 32, 33]. Revived interest in Floquet engineering, or coherent control of dressed electronic structures, has explored exotic quantum phases and symmetry-breaking in solid-state and atomic systems [34, 35, 36, 37, 38]. Engineered higher-dimensional Hamiltonians have been realized by encoding energy levels as an additional degree of freedom, or synthetic dimension [39]. Recently, ladders of $ns$ and $np$ Rydberg levels have been mapped to a synthetic dimension to probe band structure [20].
The high tunability of strong interactions among ultracold atoms make them excellent platforms for quantum simulation and studies of thermalization. Recent work has explored energy transport in the context of quantum many-body phase transitions [40, 41] and the degree of correlation between entanglement growth and diffusion [42, 43, 44]. Energy transfer through spatial redistribution of electronic states has also been realized with direct imaging [45, 46]. Alternatively, slowed transport has been observed in the presence of quantum many-body scar states [47, 48, 49, 50], and in dynamically localized systems such as the quantum kicked rotor [51, 52, 53].
In this Letter, we measure the energy transport in nearly harmonic ladders of Stark energy levels in ultracold Rydberg atoms. The energy exchange is mediated by resonant dipole-dipole interactions among the amorphous cloud of rubidium Rydberg atoms in a magneto-optical trap (MOT). The sensitivity of Rydberg atoms to external electric and magnetic fields offers some control over the parameters of the ladder, making this a promising model system.
When an atom is immersed in an electric field, the degeneracy among different angular momentum states is lifted. In non-relativistic hydrogen, the state splittings are dominated by the linear Stark effect, which leads to a harmonic spacing. While the $\ell$ degeneracy is lifted, the $m$ degeneracy is not. Thus, each rung of the ladder consists of a set of even or odd $m$ states with the number of degenerate states increasing at the center of the manifold. There are however some deviations from this in alkali atoms due to the complexity of the ion core. The quantum defects lift the $\ell$ degeneracy at zero field leading to a quadratic Stark shift at low field. Above $\ell=3$, the quantum defects are small so this quadratic shift quickly returns to a nearly linear shift as the field increases.
Including fine structure and considering only the low $m_{j}$ states in rubidium, we can visualize the manifold as clusters of six states each with $|m_{j}|=\frac{1}{2},\frac{3}{2},\frac{5}{2}$. For an initial cluster near the center of the manifold, dipole moments between clusters vary between 0 and 60 $ea_{0}$ whereas dipole moments within clusters can be much stronger, around $250$ $ea_{0}$. Clusters are separated by about 0.8 GHz and the initial cluster is about 14 MHz wide. The closest pair of states within the initial cluster are separated by about 0.4 MHz.
For a sample of atoms initially excited to a state near the middle of the Stark manifold, we might expect dipole-dipole interactions to drive the system into thermal equilibrium, with population spread uniformly across the manifold. Instead, we find that a significant fraction of the population remains in the initial state, even at long interaction times. A simple model of this system confirms this behavior, pointing to the interplay between interactions within a cluster and across rungs of the ladder as the cause of the non-thermal dynamics.
Figure 1: (a) Stark map showing the $|m_{j}|=1/2,3/2$, and 5/2 states of the $n=34$ manifold and the $36d$ states. These manifold states are organized into clusters of energy levels with nearly harmonic spacing. The initially excited manifold state is highlighted in red. During an interaction time $\leq 1\ \mu s$, resonant Förster interactions transfer population to clusters above and below the initial cluster, which are highlighted in different colors. A microwave pulse is scanned over frequency to transfer the population of each manifold state to the $36d$ state, where it can be resolved with state selective field ionization. (b) An electric field scan at a particular wavelength of the Rydberg excitation laser, showing each manifold state intersected by the horizontal dashed line in (a).
Our experiment uses 85Rb atoms that are cooled and confined in a magneto-optical trap. With the trapping laser driving the $5s\rightarrow 5p$ cycling transition, we use two additional lasers to excite Rydberg atoms. One laser operates at 776 nm to drive the $5p\rightarrow 5d$ transition and the other excites from $5d$ to states with $nf$ character at 1265 nm.
Figure 1(a) shows a Stark map for the region of interest. We must first find the desired Stark state. This is done by locking the frequency of the 1265 nm laser to an energy just below the zero-field $-\ell$ $n=34$ manifold. We then measure the number of excited Rydberg atoms as the field is increased as shown in Fig. 1(b). In this experiment we chose the state that tunes into resonance at 7.15 V/cm. For this initial state, neighboring clusters are nearly harmonically spaced and the spacings of states within each cluster is small. Once the desired state is found, the static electric field remains fixed at this value.
Upon excitation to the Stark state, atoms can immediately begin to exchange energy through dipole-dipole interactions, which spreads population to neighboring states. To study this energy transport, we must measure the distribution of population across the Stark manifold. Selective field ionization cannot resolve the signal from these closely spaced Stark states, but it can resolve the $36d$ state from those in the $n=34$ manifold. We therefore use a pulsed microwave field as shown in Fig. 1(a) to drive population from the manifold states to the $d$ state. As the microwave frequency is scanned we measure the $d$ state population to reveal the spectrum of excited Stark states as shown in Fig. 2.
Figure 2: Field ionization signal of the $36d$ state as a function of microwave frequency, shown at different delays. Each shaded region corresponds to a cluster of manifold energy levels in Fig. 1(a). Each peak within a shaded region is due to a different $36d|m_{j}|$ sublevel. All five possible peaks are most easily visible in central, red-shaded region corresponding to the initially excited state. With increasing delay, Förster interactions transfer population from the initial cluster to adjacent clusters. For 0 $\mu s$, the end of the Rydberg excitation pulse coincides with the start of the microwave pulse.
Each panel in Fig. 2 shows the microwave spectrum for a particular time delay between the laser excitation pulse and the microwave probe pulse, which is the time during which atoms can interact and exchange energy. Both pulses have a width of 50 ns and a delay of 0 ns marks the time at which the trailing edge of the laser pulse is aligned with the rising edge of the microwave pulse. At 7.15 V/cm, the $36d$ state is split into five states; $j=\frac{5}{2},\ |m_{j}|=\frac{1}{2},\ \frac{3}{2},\ \frac{5}{2}$, and $j=\frac{3}{2},\ |m_{j}|=\frac{1}{2},\ \frac{3}{2}$. The range of energy between these five states is slightly smaller that the separation between neighboring manifold states. We therefore expect the microwave spectrum to show five peaks for a single Stark state cluster. The colored bands in Figs. 1 and 2 mark the range of microwave resonances associated with a single cluster of Stark states. To find the population in each Stark cluster, we integrate the signal in each band. The resulting energy transport due to dipole-dipole interaction is shown in Fig. 3 where the color and dashing of each curve matches that of Figs. 1 and 2.
[…]
At an interaction time of $-100$ ns there should be no signal in the $d$ state since the microwave probe pulse turns off before the Rydberg excitation pulse turns on. In fact, we see signal in the $36d$ state due to the finite turnoff time of the microwave pulse. As the delay is increased to 0 ns, we see the overall signal rise as well as population in neighboring Stark states. Beyond 0 ns, the interaction continues to spread population to neighboring manifold clusters although this spread is rather slow, leaving the bulk of the population in the initial Stark state. In Fig. 3(a) we also see an asymmetry between clusters above (dashed lines) and below (solid lines) the initial state. This is largely due to the variation in the microwave coupling strength between manifold states and the $36d$ state.
To measure the microwave coupling strengths, we switched to exciting the $36d$ state [54]. We measured, using selective field ionization, the population in the $n=34$ manifold whilst scanning microwave frequency. By separately exciting the five $36d$ states, we produced a map of the coupling strength across the manifold for each $36d$ state, which were weighted to match data shown for the initially excited state in Fig. 2. Using the same integration bins as shown in Fig. 2, we found the overall weighting factors, which were used to scale the data in Fig. 3(a). The result of this scaling, shown in Fig. 3(b), reveals a more symmetric population of clusters above and below the initially excited Stark state.
Ideally, we would quench the system into resonance. However, since the dipole-dipole interactions are resonant at a wide range of electric fields, we cannot excite at a non-resonant field. Thus, interactions occur during the 50 ns laser excitation pulse and some of the energy transport visible before 0 ns in Fig. 3 is possibly due to direct excitation of pair states [55]. However, after 0 ns the initial state population continues to decrease while the populations of other clusters increase, a clear signature of dipole-dipole interactions. In fact, Seiler et al. determined that resonant Förster interactions among manifold states were required to model the decay of Rydberg atom population in an electrostatic trap [56].
[…]
Using the calibrated density we model the manifold system by numerically solve the Schrödinger equation using the two-body Hamiltonian
$\sum_{a}\sum_{i}\hat{\sigma}_{aa}^{i}E_{a}+\sum_{a\neq b}\sum_{c\neq d}\sum_{i\neq j}\left(\hat{\sigma}_{ab}^{i}\hat{\sigma}_{cd}^{j}+\mathrm{H.c.}\right)\frac{\mu_{ab}\mu_{cd}}{R_{ij}^{3}},$ (1)
where $\hat{\sigma}^{i}_{ab}$ is an operator that takes the $i^{th}$ atom from Stark level $a$ to Stark level $b$ and $R_{ij}$ is the distance between the $i^{th}$ and $j^{th}$ atoms. The transition dipole moments $\mu_{ab}$, which connect Stark level $a$ to $b$, are calculated using the Numerov method [58]. The first term gives the energy of each level and the second term represents the two-body Förster interactions. When $c=a$ and $d=b$, the second term yields hopping interactions.
Our simulations assume that the atoms are motionless and we ignore the angular dependence of the interactions. We randomly place dozens of atoms in a spherical volume at our measured density. Including enough energy levels to faithfully represent the manifold necessarily limits the number of atoms we can include in the Schrödinger equation. We iterate over each atom in the volume and choose its three nearest neighbors and then run the simulation with those four atoms so that we include the possibility of “secondary” interactions. For example, two pairs of atoms in the initial cluster could exchange energy, resulting in a pair of atoms in other energy clusters that are now able to interact. We repeat this process for multiple volumes and average so that the final results include a few hundred simulations.
With only four atoms in the model, the simulated dynamics will be slower than in the experiment [50]. Since we also do not model direct excitation of pair states, our simulated evolution is not suited for quantitative comparison to the experiment. However, we can look at longer simulated times, when the populations of each energy cluster have plateaued.
Figure 4: Fraction of atoms as a function of energy for a simple harmonic ladder (black) and a more complete model of the Stark manifold (red). The fraction of atoms in each energy level is mapped to the area of a gaussian of width 0.05 GHz. The small anharmonicity of the full model is visible as a progressive misalignment of the peaks.
As a starting point for understanding the spread of energy in our system, we consider a perfectly harmonic ladder of nine energy levels with four atoms initially in the middle level. We set the separation between energy levels to be 0.8 GHz, the mean separation between clusters in the actual experiment. The $\mu_{ab}$ are randomly drawn from a box distribution centered on the mean between-cluster dipole moments. Figure 4 shows the distribution of energy, averaged over the last five microseconds of a 100-microsecond simulation. The system has thermalized, with a relatively even energy distribution. When we run this model with five or six atoms and more energy levels, we find that the system still thermalizes.
The predictions of the harmonic ladder disagree strongly with the experimental result. We can attempt to improve the model by adding some anharmonicity. Moving outward from the initial cluster used in the experiment, the mean energy of each cluster shifts by about 0.1% of the cluster spacing. With this small anharmonicity, we find that the results do not significantly change and the system still thermalizes.
We further attempt to make the model more realistic by adding clusters of energy levels. We include nine clusters of four $|m_{j}|=\frac{1}{2}$ and $\frac{3}{2}$ energy levels, which are taken from the calculated Stark energy levels from Fig. 1 along with the calculated $\mu_{ab}$. We include only four-atom states whose detuning from the initial state is less than the typical energy spacing between clusters, resulting in about 125 000125000125\,000125 000 states.
While this model still ignores many energy levels with $|m_{j}|>\frac{3}{2}$, the simulation results shown in Fig. 4 are in qualitative agreement with the data shown in Fig. 2. The model shows a similar degree of symmetric energy spreading to more distant clusters, while the bulk of the population remains in the initial cluster. The dynamics are strikingly simple given the complex structure of the energy levels and the couplings $\mu_{ab}$.
Some of this simplicity may be attributed to our choice of initial cluster. As seen in Fig. 1, the initial cluster is so tightly spaced that the individual energy levels are not resolved. The bandwidth of our excitation laser is broad enough to excite a superposition of these energy levels. With so many possible initial pair states, there are dipole-dipole interactions with small detuning to every final pair of clusters. Interference among the many pathways from initial state to final state may play a role in the failure of this system to thermalize.
Our results indicate that the Rydberg Stark manifold could offer a promising platform for studying quantum dynamics with a rich and complex energy level structure. Experiments studying dipolar energy exchange in atomic systems usually start with a simpler set of energy levels. The many closely-spaced energy levels in the manifold are more similar to solid-state systems, offering prospects for quantum emulation. While the energy levels discussed in this experiment are arranged into nearly harmonic clusters, the addition of a static magnetic field could either broaden those clusters or cause them to overlap. For $|m_{j}|>\frac{5}{2}$, there are also nearly harmonic ladders that fall between the ladder studied here.
Interactions beyond two-body energy exchange are also possible in the manifold. In fact, nearly every two-body interaction studied here is somewhat detuned from resonance. Since there are many energy-level spacings available within the clusters, the addition of a third atom can bring the interaction closer to resonance. While these three-body interactions scale as $R^{-6}$, with their smaller detuning we expect that they contribute significantly at higher densities. A study of the density dependence of energy transport in the manifold may be able to detect their effect.
Acknowledgements.
This work was supported by the National Science Foundation under Grants No. 2011583 and No. 2011610, and S.E.S. is supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2334429.This work used the Delta system at the National Center for Supercomputing Applications through allocation PHY230142 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.